OpenStudy (anonymous):

−8≤x≤8 how to put that in interval notation?

OpenStudy (anonymous):

?

OpenStudy (psymon):

If you have the greater then or less than or equal to signs, you'll use brackets when you do interval notation. When they are simply less than, greater than, you use parenthesis. If you ever have infinity or negative infinity, you always use parenthesis. As for how to put it in interval notation, you start from the most negative number you have, which is -8 in this case. Since this has a less than or equal to sign, I would put [-8, After the first number you put comma, then the most positive number you have, which is 8. Since this also has one of the equal to signs, this would also be in brackets. SO you would have [-8,8] As your answer. For a few other examples, let's say I had these: \[x < 8\] This becomes \[(-\infty, 8)\] In parenthesis because it was simply a less than sign, no equal. Another example \[x \ge -5\] this time I would have \[[-5,\infty)\] bracket with the -5 because it was an or equal to sign, off to infinity where infinity always has a parenthesis. As for a last example, this is if you had multiple intervals: \[x \le 0, x \ge 5\] If you had a situation like this, you would have to put two intervals with a union symbol in between. So I start with negative infinity up to 0 where the 0 will have a bracket \[(-\infty,0]\cup \] That U shaped sign means union. It basically represents a gap in your interval. o now after that union sign, ill put the other part of the, which start at 5 with a bracket and go to infinity. So the whole interval notation would be: \[(-\infty,0]\cup[5,\infty)\] I know its a long explanation, but I wanted to show ya the other examples, considering you may have more problems with these.

OpenStudy (anonymous):

(−∞,0]∪[5,∞) so thats the answer? the reason im asking is because someone on here told me a different answer which was just [-8,8] and i thought it didnt look right to me

OpenStudy (psymon):

No, I put [-8,8] up top xD But I was just putting down all the possible situations you could run into x_x sorry for the confusion, but yes, [-8,8]

OpenStudy (anonymous):

is that a soultion set or just solution?

OpenStudy (psymon):

Well, technically its a set. But either way, it's what you would be writing down if you needed your example in interval notation. Yeah, sorry if all the other stuff is confusing or meaningless to ya.

OpenStudy (anonymous):

no i understand it thanks! Ill rate you best response.

OpenStudy (psymon):

Alright, thanks./ Good luck ^_^